Shift summary for July 31,1999

S. Elizabeth Turner (eturner@kudzu)
Sun, 1 Aug 1999 00:02:13 -0500

Shift Summary for evening shift, July 31, 1999

On Shift: BobT, SET

Hours of beam available at all intensities: 5
Hours of beam available at nominal intensity: 5
Hours of E832 data (special runs not included): 5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: TeV quench,TECAR problems

========================================================================
Runs ended this shift:
none
========================================================================
Tapes brought to FCC this shift:
none
========================================================================
Accesses during shift:
none
========================================================================
Overview of shift:

Smooth 832 datataking until TeV quench at 20:20.
Waited for beam to return for remainder of shift.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14077 15:30 Note 2 red SDAQ buttons: TRD gas -- fixed by Sasha
G. later in shift; EPICURE -- NMU1 slightly out
of bounds (off and on all shift), Rick F. notified.
17:30 Called MCR about SEM of 3.8E12, they say this is
the best they can do w/o going to 2 turns.
223 18:20 TeV quench.
21:00 DN: 1 new bottle of CO2 on line.
22:55 DC2X tripped, Hogan talked us through the reset process
23:00 TECAR problems, no estimate for beam.
23:50 Marty from MCR stops by to explain that TECAR is
'Tevatron Excitation Control and Ramp' program.
Not clear if it's a software or hardware problem.
The guy who owns the program is out of town, the
second guy is en route. No estimate.
23:55 Marty calls to tell us everything is ok now. Est
10 min.
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

- Keep taking E832 data: no muon runs, no magnet polarity changes.